# Sorting-based FPGA Sliding Window Aggregation Engine without off-chip Memories

### v2

@@ 3.2. Sliding window aggregation engine
A sorter is needed right before the aggregation engine to sort the window based on the $group$ IDs, as well as the keys. Note that sorting by the keys as well is not always required, such as when the operator is commutative as with summation. Since 4K elements, for example, are considered moderately large as SWAG window sizes (Geethakumari and Sourdis, 2023), the sorting can entirely be done on-the-fly instead of indexing larger memories. FPGA-based high-throughput sorters are thoroughly studied in literature (Saitoh et al., 2018; Zuluaga et al., 2016; Qiao et al., 2022; Song et al., 2016).
Steps (c) and (d)
@@
These steps remain similar as with the design of section 3.1, but with one of the three simultaneous cumulative sums (for keys, valid bits and last bits) decoupled from the others, in order to happen first. Hence, it needs two prefix scan-like structures. The preceding prefix scan is on the valid bits, so that the index of each tuple within the group is known first. The idea is to be able to do selection-based queries such as the median value, just by comparing this index with the group cardinality that is already-appended to all tuples by the modified sorter (section 3.3). In this work, we focus on the min./med./max. operator, which leaves the first, middle, and last tuples of each group (corresponding to the min./med./max.) as a valid aggregate result, as opposed to a single tuple for the aggregated summary of the group-by engine. If queries such as the median are not needed, the group-by engine of section 3.1 is sufficient to perform its simpler operators in a sliding-window context.
Figure 6. SWAG numerical example: finding the minimum, median and maximum value per group per sliding window.
Steps (e) and (f)
@@
The last two steps reduce back to a PRRA architecture (Papaphilippou et al., 2023), as the remaining functionality only relates to the round-robin compaction. The final index calculation is done by the last prefix scan in the pipeline, while the first one is now responsible for all computation-related functionality including the other two cumulative sums. See the numerical example of figure 6, which has similar semantics as figure 4. It demonstrates the min./med./max. selection, as well as the expected positions of the results in the output after the abstracted steps (e) and (f) as a PRRA.
3.3. Sorter with group-cardinality calculation
@@
Additional logic is needed when supporting non-incremental aggregation functions such as when finding the median. The calculation of the median value is not trivial without knowing the size of the data for which it is applied. This requires a modification inside the preceding sorter to be able to append the median-related information, which is the group cardinality alongside the data passed on to the aggregation engine. This functionality can conveniently be integrated into the sorter, as the sorter is provided with the entirety of the windows before flushing its sorted output to the aggregation engine.
@@
The ability to involve median calculation within a sorter depends on the sorter architecture (Guo and Luk, 2001). Our proposed solution is based on a sorter that uses the linear sorter as a building block (Papaphilippou et al., 2020). The linear sorter is the hardware equivalent of insertion sort, as it is able to insert a new value into an already sorted sublist (Koch and Torresen, 2011). Each sorting cell, as depicted in figure 7 (center), has a comparator that compares its tuple with the incoming one. According to the result of the comparison and the one of its neighbour (“left smaller”), it either adopts the new value, propagates the left value, or remains as is.
Figure 7. Adding group-cardinality support to the linear sorter.
@@
The baseline sorter of our design is able to provide high-throughput (multiple tuples per cycle) while supporting moderately large window sizes. The open-source sorter of (Papaphilippou et al., 2020) is specifically selected because it combines linear sorters with a merge tree (PMT (Song et al., 2016)) for high-throughput. This combination features special adaptations to be able to work efficiently in many-leaf (input list) merge mode using the same hardware.
Figure 8. Data movement inside the proposed SWAG engine, according to the window size (WS) and internal size ($k$) of the merge-capable sorter.
@@
Figure 9. Adding group-cardinality support to the sorter.
@@
An additional modification is needed to propagate the correct group cardinalities, as summarised in figure 9. The parallel merge tree (PMT) needs to sum the cardinalities calculated by the linear sorters. Each FLiMS merger (Papaphilippou et al., 2022) of the PMT uses a network of compare-and-swap (CAS) units, which are sorters of 2-elements. The modification here is that each CAS unit uses the cond1 and cond2 conditions to determine how to update two incoming cardinalities every time they share the same group ID. The tuples carry two extra bits inside the mergers, one for denoting if the updating of their cardinalities has already happened, and one for their original source (lists A or B according to figure 9). Cond1 (to sum both cardinalities) is true whenever the cardinality update has not already happened to any of the 2 inputs, and the source is different. Cond2 (to adopt the other cardinality) is true whenever the cardinality update has happened to the other tuple, and the source is either different or the same. A similar arrangement is done for the selector stage units, as they are based on the same comparison, but they also remember any merged cardinalities from the elements of the last comparison. The latter is needed to support the case when a set of same group IDs is only coming from one of the lists A or B, such as when the ones from the other list have already been extracted.
The idea of a proof for this algorithm relates to the fact that the tuples sharing the same group ID (most significant bits) will appear consecutively in the output of the FLiMS merger, since its output is always sorted. As the order of those “duplicates” in regard to their group ID will also be correct in regard to their other bits, this means the required pairwise comparisons take place to establish this order, which are enough to broadcast such information, as with spanning trees (Touzene and Day, 2004).
4. Evaluation
This section evaluates the implementations of the group-by aggregate and SWAG engines. The studied pipelines accept a data stream consisting of $(group,key)$ pairs. The engines are highly parameterisable, but the evaluation mostly focuses on the following configuration: each tuple is 64-bit wide, 32-bit for each of the two fields, a pipeline throughput $P$ of 4 tuples per cycle, and a sorter capacity $k$ of 128 (supporting window sizes of up to $\textit{WS}=16K$ for the SWAG engine). The WS and WA values are configured at run time.
@@ 4. Evaluation
All presented designs are written in Verilog [^1] Available: (open source repository will be released after peer-review). The entities such as the compare-and-swap units for the sorter’s merger cells or the “$n$” entities of the prefix scan are parameterisable, such as with custom data widths and functionality, and are written by hand. The structures that use those entities including the reverse butterfly network and the parallel merge tree in the sorter are generated by scripts. These Verilog-generating scripts are written in Python to ease the inference of more complex or repetitive entity topologies for arbitrary $P$ values, as well as $k$ values for the sorter.
4.1. SWAG engine throughput and latency
[…]
Figure 10. Final input rate influenced by backpressure, $P=4$.
[…]
Figure 11. Simulated latency, $P=4$, $k=128$, $f_{\textit{target}}=125$ MHz
Figure 11 presents a similar experiment that measures the latency between the time the first tuple enters into the engine and the time there is a tuple in its output. As it can be observed, the WS does not influence the latency up to $\textit{WS}=128$, because the linear sorters in the sorter have a fixed latency ($k/P$) for an input to be extracted. For $\textit{WS}>k$ the latency scales linearly with WS, because the sorted windows are merged with a fixed rate $P$, and the merge phase is compulsory. WA does not impact the latency.
[…]
4.2. System-level evaluation
@@
The resulting engines are evaluated on the Ultra96 development board with the Zynq UltraScale+ ZU3EG device, an MPSoC that runs Linux on its ARM A53 cores. This is an embedded platform, showcasing the resource-efficiency and adaptability of the proposed SWAG engine.
@@
The Verilog code is implemented on the FPGA using AMD Vivado. The 2GB memory of the board is divided in two halves by the device tree, the kernel-mapped part, which is optimised for cached access to the CPU, and the non-kernel part, which is appropriate for FPGA use. The operating frequency of all designs is 250 MHz, with an internal frequency of 125 MHz, resulting in a 256-bit/cycle datapath. The critical path is coming from the sorter, which highlights the benefit of being able to alleviate a lower operating frequency of FPGAs with high-throughput processing. Still, 250 MHz is relatively close to saturating the memory bandwidth for this platform (Papaphilippou et al., 2020). Figure 12 visualises this experimental setup.
Figure 12. Experimental setup on the MPSoC platform.
Figure 13. SWAG engine speedup over one A53 ARM core for 3 operator types.
@@
The data first pass through the high-throughput sorter, and its output is directed towards the proposed aggregation engines. The engine is either the group-by-aggregate version or the corresponding part of the SWAG engine.
@@
The core uses a direct memory access (DMA) engine to transfer the data from the main memory to the sorter. The engine configuration such as the operator selection is managed by writing to a few memory-mapped registers. As soon as the data arrives to the sorter, they are immediately being sorted into chunks of up to $k$ size. The destination of those sorted chunks are either a temporary block memory (BRAM), the main memory, or the aggregation engine. Since the proposed solutions use a merge sorter, it is a recursive operation, and the engine only reads the sorter’s output on the last merging pass, hence the multiplexer.
@@
The sorter does not currently need to output data to the main memory for the SWAG version, however, as the $k^{2}$-sized windows are deemed enough. Hence, the DRAM here is part of the evaluation platform and is not involved or needed in the calculation of the SWAG queries. The handling of the data by the sorter is preceded by a simple buffering mechanism that provides the studied windows in their entirety, as described in section 3.2.
The designs are compared against highly-optimised C++ implementations. The algorithms are relatively simple and they rely heavily on well-studied functions and structures from the C++ Standard Library. The GCC compilation flags are -O3 -march=native. These only target a single of the four available cores to avoid introducing more design variables. Due to Amdahl’s law and other inefficiencies coming from multi-processing, we can consider that the obtained FPGA speedups are at most roughly four times higher than multi-threaded implementations.
4.3. Group-by aggregation (preliminary experiment)
The proposed group-by-aggregate engine is adapted as a building block in the novel SWAG engine. This part of the evaluation preliminarily explores it as a standalone accelerator. The supported operations per $group$ ID in this implementation are: minimum, maximum, sum, count and distinct count.
@@ 4.3. Group-by aggregation (preliminary experiment)
The C++ baseline is equivalent to the functionality of the design, i.e. first sorting the data and then performing the aggregation serially. The sorting function is std::sort, which is a well regarded sorting baseline for generalisable use, although not the absolute fastest. The achieved speedup for $2^{16}$ tuples over the embedded processor varies between 22 to 28 according to the input distribution. This variation relates to the number of rows in the output. For instance, if all tuples have the same group ID, then the latency of interacting with the FPGA would be comparatively higher, as only a single tuple is produced as an output.
@@
This speedup mostly comes from the sorter, as the studied CPU is already efficient for linearly skimming through a sorted stream while aggregating, also given its advantageous memory interfacing and speed. The speedup behaviour, such as for arbitrarily long streams, is inline with the sorter (Papaphilippou et al., 2020), hence not elaborating with comparison details. The advantages in resource utilisation are described in section 4.5, and the discussion of this standalone functionality in section 5.
4.4. Sliding window aggregation (SWAG)
The proposed SWAG engine is implemented as an AXI peripheral, and is used as a co-processor in our Linux-running evaluation environment. The supported operators are the ones of the latter plus the min./med./max. operator. It provides these aggregation results in the form of up to 3 tuples per group representing the minimum, median and maximum with the corresponding group ID. This is achieved by the proposed approach as introduced in section 3.2 that only leaves this subset to pass as valid after the first prefix scan in the pipeline. Since this is based on filtering to achieve aggregation, when there are any overlappings, such as with a singleton group whose minimum tuple is the maximum, such extracted tuples are considered common and the number of output tuples never surpasses the input tuples. In all other aggregation queries, the last tuple of each group is the one that gets updates to carry the aggregation summaries, as with the unmodified group-by aggregate engine of section 3.1.
@@ 4.4. Sliding window aggregation (SWAG)
The query selection is sum, distinct count and min./med./max. The logic behind this selection is to select a variation of algorithms that benefit the CPU and the FPGA at different degrees. For instance, the sum operator is incremental, and a pointer-based solution with software can potentially have an advantage over the FPGA solution, as it does not have to read entire windows for smaller window updates. On the other hand, min./med./max. is less trivial or rewarding to calculate incrementally on the CPU. The baseline software is based on C++ std::set and std::map to hold and update per-group information (notably similar performance to their hash-based unordered counterparts, not shown), and std::nth_element() to calculate the medians.
@@
As demonstrated in figure 13, the FPGA advantage of the solution is different for the three operator types. There is a clear influence of the throughput behaviour, as was seen in the simulated SWAG engine of figure 10, since after $\textit{WS}>k=128$, there is the internal second pass to support the larger windows that halves the throughput. Still, for the majority of operators, WS and WA combinations, the three operators provide a speedup of one to two orders of magnitude. If we ignore the dip after the internal pass is used ($\textit{WS}<128$), we can see the upward trend in the figures with respect to the window size, which can be explained by the thoughtful use of the on-chip memory and longer memory transfers, as opposed to the CPU’s cache limitations. An exception is the combination of a high WS and low WA for the sum operator where the FPGA experiences a speedup lower than 1x. This comes from the software advantage of being able to efficiently do incremental updates for such a simple operator as sum, though according to a use case such as with 100% unique group IDs, WS=128, WA=64 the speedup is 164x (not shown).
@@
The studied operators also behave differently under different input distributions at least when compared to the software equivalent. The input of sum and distinct count operators consists of 50% unique group IDs per group, which means every window has around WS/2 unique group IDs. The variations when changing from 100% unique to near 0% unique (single group ID) were relatively limited for these two operators, especially with regard to the FPGA performance, hence only the 50% version is shown to conserve space. For example, for WS=16384, WA=8192 the speedup on the distinct count operator is 317x, 420x and 476x for 100%, 50%, and single unique groups respectively. However, this is not the case for our main target min./med./max., hence the combined rightmost plot of figure 13. The output size varies according to the distribution, ranging from 1 to 3 tuples per group ID, according to the overlappings. If we select WS=16384, WA=64, the corresponding speedups for min./med./max. change to 59x, 239x and 472x. This is expected as the increased localities from having to operate on fewer elements can benefit software implementations when comparing the overheads of using a co-processor, while noting that the speedup is still impressive.
4.5. Resource utilisation
Scalability (synthesis-based)
The proposed aggregation engines are scalable to arbitrary degrees of parallelism $P$ to be able to saturate the bandwidth for current and future systems with a wide datapath. The total amount of data that can be produced or consumed by the engines per cycle is $P$ times data width. An exploration of the impact of $P$ and the data width on the resources is conducted with out-of-context synthesis results. The tuple (data) width is halved and used by the group ID and key fields for simplicity. The synthesis is done with Yosys (Wolf, 2016) for a more generic but modern 6LUT architecture, which can serve as high-level guidelines for the resource utilisation expectations in future applications.
Figure 14 summarises the results of this exploration for the group-by engine (upper) and the SWAG engine (lower), without the accompanying sorters and sliding window buffering. Since the first uses the unmodified open-source sorter, and the latter an adaptation of it, the readers can also refer to its original description (Papaphilippou et al., 2020) or an alternative implementation.
Figure 14. Synthesis results when varying tuple width and $P$ (sorter excluding).
For the ranges of interest, the registers and LUTs scale almost linearly with $P$ and the tuple width. The fastest growing component is the reverse butterfly network, with a complexity of $\Theta(P\log_{2}(P))$ switches (similar for the prefix scans), though the slight upward trend is not observed in these ranges yet. The register utilisation is always more for the SWAG engine than the group-by engine, and this can be explained by the decoupling of the simultaneous prefix scan part into two consecutive ones. This has lengthened the pipeline length, and thus, register use, as the presented implementations use a register per pipeline stage. Our use case has yielded a low enough register utilisation for our purposes, but future work can also focus on retiming to considerably reduce the register utilisation when necessary.
The LUT utilisation for the group-by engine and SWAG engine are similar, but there is variation between their corresponding configurations, as sometimes SWAG is more lightweight. Although the additional pipeline stages and the additional support of min./med./max. has an overhead in utilisation, many of the row/cardinality counters are narrower than the corresponding of the group-by engine. This is since the scope of the window size (WS) for the SWAG engine are limited to $k^{2}$ (here equal to $128^{2}$). The group-by engine uses counters of key-width, which also explains the higher impact from the tuple width in the upper plots of figure 14.
Overall, $P$ seems more influential to the LUT and register use than the tuple width, as it has the potential to scale to more orders of magnitude for the studied ranges. Even the highest points here ($P=64$, tuple width 256 bits) are feasible with today’s datacenter FPGAs, though any complications would first relate on how easy is to establish a sorted stream with such a high aggregate throughput.
Implementation

### v3

@@ III-B Design
a sorter is needed right before the aggregation engine to sort the window based on the groupgroup IDs, as well as the keys. Note that sorting by the keys as well is not always required, such as when the operator is commutative as with summation. Since 4K elements, for example, are considered moderately large as SWAG window sizes [7], the sorting can entirely be done on-the-fly instead of indexing larger memories. FPGA-based high-throughput sorters are thoroughly studied in literature [8, 9, 10, 11].
Step (c) Prefix-scans
this step remains similar as with the design of section II-B, but with one of the three simultaneous cumulative sums (for keys, valid bits and last bits) decoupled from the others, in order to happen first. Therefore, it now requires two prefix scan-like structures. The preceding prefix scan is on the valid bits, so that the index of each tuple within the group is known first.
Step (d) Rollover and summarisation
@@
the idea is to be able to do selection-based queries such as the median value, just by comparing this index with the group cardinality that is already-appended to all tuples by the modified sorter (section III-D). One target operator is min./med./max., which leaves the first, middle, and last tuples of each group (corresponding to the min./med./max.) as a valid aggregate result, as opposed to a single tuple for the aggregated summary of the group-by engine.
Steps (e) 2nd Prefix-scan and (f) Stream compaction
@@
the last two steps reduce back to a PRRA architecture [4], as the remaining functionality only relates to the round-robin compaction. The final index calculation is done by the last prefix scan in the pipeline, while the first one is now responsible for all computation-related functionality including the other two cumulative sums.
III-C SWAG numerical example
Figure 5: Colour-coded example on finding the min./med./max. per group.
The numerical example of figure 5 demonstrates the min./med./max. selection, and the expected positions of the results. The idea for the sorted-supplied cardinality information is to be able to perform selection just by comparing the cumulative group cardinality (position of a tuple inside its sorted group) with the pre-calculated one (section III-D).
III-D Sorter with group-cardinality calculation
@@
Additional logic is needed when supporting non-associative aggregation functions such as when finding the median. The calculation of the median value is not trivial without knowing the size of the data for which it is applied. This requires a modification inside the preceding sorter to be able to append the median-related information, which is the group cardinality alongside the data passed on to the aggregation engine. This functionality can conveniently be integrated into the sorter, since the sorter is provided with the entirety of the windows before flushing its sorted output to the rest of the engine.
@@
The ability to involve median calculation within a sorter depends on the sorter architecture [14]. Our proposed solution is based on a sorter that uses the linear sorter as a building block [15]. The linear sorter is the hardware equivalent of insertion sort, and is able to insert a new value into an already sorted sublist [16]. Each sorting cell, as depicted in figure 6 (center), has a comparator that compares its tuple with the incoming one. According to the result of the comparison and the one of its neighbour (“left smaller”), it either adopts the new value, propagates the left value, or remains as is.
Figure 6: Adding group-cardinality support to the linear sorter.
@@
The baseline sorter of our design is able to provide high-throughput (multiple tuples per cycle) while supporting moderately large window sizes. The open-source sorter of [15] is specifically selected because it combines linear sorters with a merge tree (PMT [11]) for high-throughput. A special feature of this combination is that it is able to work efficiently in many-leaf (input list) merge mode using the same hardware, and this also applies to our new logic for resource efficiency.
Figure 7: Data movement inside the proposed SWAG engine, according to the window size (WS) and internal size ($k$) of the merge-capable sorter.
@@
Figure 8: Adding group-cardinality support to the high-throughput sorter.
@@
An additional modification is needed to propagate the correct group cardinalities, as summarised in figure 8. The parallel merge tree (PMT) needs to sum the cardinalities calculated by the linear sorters. Each FLiMS merger [17] of the PMT uses a network of compare-and-swap (CAS) units, which are sorters of 2-elements. The modification here is that each CAS unit uses the cond1 and cond2 conditions to determine how to update two incoming cardinalities every time they share the same group ID. The tuples carry two extra bits inside the mergers, one for denoting if the updating of their cardinalities has already happened, and one for their original source (lists A or B according to figure 8). Cond1 (to sum both cardinalities) is true whenever the cardinality update has not already happened to any of the 2 inputs, and the source is different. Cond2 (to adopt the other cardinality) is true whenever the cardinality update has happened to the other tuple, and the source is either different or the same. A similar arrangement is done for the selector stage units, as they are based on the same comparison, but they also remember any merged cardinalities from the elements of the last comparison. The latter is needed to support the case when a set of same group IDs is only coming from one of the lists A or B.
IV Evaluation
This section evaluates the implementations of the group-by aggregate and SWAG engines as accelerators. All implementations are accompanied by a sorter/ cardinality counter of capacity $k$ of 128 elements, supporting arbitrary input sizes for the group-by engine by being applied recursively, and window sizes of up to $\textit{WS}=16K$ for the SWAG engine. The operator type is configured at run-time. The implemented engines support multiple operators also configurable at run-time, as summarised in table I.
IV-A Experimental methodology
The engines are evaluated as co-processors in a Linux-running environment. They are implemented as AXI peripherals, and the codebase is mainly written in Verilog [^1] Source code (base): https://philippos.info/groupby/
Source code (SWAG): available soon
@@ IV-A Experimental methodology
. The entities such as the compare-and-swap units for the sorter’s merger cells or the entities of the prefix scans are parameterisable, such as with custom data widths and functionality, and are written by hand. The structures that use those entities including the reverse butterfly network and the parallel merge tree in the sorter are generated by scripts. These are written in Python to easily infer such complex or repetitive entity topologies for arbitrary $P$ and $k$ (i.e. sorter size) values.
@@
The resulting accelerators are evaluated on the Ultra96 development board with the Zynq UltraScale+ ZU3EG device, an MPSoC that runs Linux on its ARM A53 cores. This is an embedded platform, showcasing the resource-efficiency and adaptability of the proposed engines, while being scalable for higher-end platforms as well.
@@
The 2GB memory of the board is divided in two halves by the kernel; the kernel-mapped part, which is optimised for cached access to the CPU, and the non-kernel part, which is left for the FPGA. The target operating frequency of the AXI-related IPs is 250 MHz, with an internal frequency of 125 MHz for the adapted sorter and engines, complying with the sorter’s original design. This results in a 256-bit/cycle datapath for the engine from an 128-bit-wide HP port. This is relatively close to saturating the memory bandwidth of this platform [18]. Figure 9 visualises this experimental setup.
Figure 9: Experimental setup on the MPSoC platform.
@@
The data first pass through the high-throughput sorter, and its output is directed towards the rest of the aggregation engine. The engine is either the group-by-aggregate version or the corresponding part of the SWAG engine. For the first, the sorter is the unmodified open-source IP from [15], whereas the latter combines the group cardinality functionality from section III-D.
@@
The core uses a direct memory access (DMA) engine to transfer the data from the main memory to the sorter-cardinality counter. The engine configuration such as function_select is managed by writing to a few memory-mapped registers. As soon as the data arrives at the sorter, they are immediately being sorted into chunks of up to $k$ size. The destination of those sorted chunks are either a temporary block memory (BRAM), the main memory, or the aggregation engine. Since the proposed solutions use a merge sorter, it is a recursive operation, and the engine only reads the sorter’s output on the last merging pass, hence the multiplexer.
@@
The sorter-cardinality counter does not currently need to output data to the main memory for the SWAG version, however, since the $k^{2}$-sized windows are deemed enough. Thus, the DRAM here is part of the evaluation platform and is not involved or needed in the calculation of the SWAG queries. The handling of the data by the sorter is preceded by a simple buffering mechanism that provides the studied windows in their entirety, as described in section III-B.
Their performance is compared against highly-optimised software, including reference implementations from IBM for SWAG [19]. The custom operators remain relatively simple, as they rely heavily on well-regarded functions and structures from C++’s standard template library (STL). The baseline software uses std::set and std::map to hold and update per-group information (notably similar performance to their hash-based unordered counterparts, not shown), and std::nth_element() to calculate the medians. The GCC compilation flags are -O3 -march=native.
IV-B Experimental results – Group-by aggregation
This part of the evaluation explores the base pipeline as a standalone add-on to the existing unmodified small-scale sorter [15]. The supported operations per groupgroup ID in this implementation are: minimum, maximum, sum, count and distinct count. The engine accepts a data stream consisting of $(group,key)$ pairs. Each input tuple is 64-bit wide, 32-bit for each of the two fields. The pipeline throughput $P$ is 4 tuples per cycle.
@@ IV-B Experimental results – Group-by aggregation
The C++ baseline is equivalent to the functionality of the design, i.e. first sorting the data and then performing the aggregation serially. The sorting function is std::sort, which is a high performing sorting baseline for general use, although not the absolute fastest [17]. The achieved speedup for $2^{16}$ tuples over the embedded processor varies between 22 to 28 according to the input distribution. This variation relates to the number of rows in the output. For instance, if all tuples have the same group ID, then the latency of interacting with the FPGA would be comparatively higher. This is because only a single tuple would be produced as an output.
@@
This speedup mostly comes from the sorter, as the studied CPU is already efficient for linearly skimming through a sorted stream while aggregating, also given its advantageous memory interfacing and speed. The speedup behaviour, such as for arbitrarily long streams, is inline with the sorter [15], hence not elaborating with comparisons for standalone group-by.
IV-C Experimental results – SWAG engine, without groups
The SWAG engine is initially evaluated without group IDs, and a single stream of 32-bit integers is processed with high throughput. In this case, in order to saturate the 256-bit datapath, the generated aggregation pipeline and sorter-cardinality counter are for $P=8$, and without groupgroup fields. As with the operator type, the WS and WA values are also configured at runtime.
This engine is compared against IBM’s reference implementations of multiple aggregation algorithms on the platform’s CPU [19] after porting it to ARM. These include the nbfinger, amta, two_stacks_lite, chunked_two_stacks_lite, daba_lite variations using the run_bulk_evict_insert.py experiment setup. The tested operations are the commonly supported sum and min. or max. for data in-order, though only sum is shown for brevity, as the performance remains relatively similar for these operators. Each kept datapoint is the minimum time from all algorithms, so that we compare with the best available software performance per window size (WS) and window advance (WA) combination.
Figure 10: SWAG, without groups: engine speedup over one A53 ARM core.
Figure 11: SWAG, with groups: engine speedup over one A53 ARM core for 3 operator types.
Figure 10 summarises this exploration along a custom made distinct count routine, which is representative of more advanced queries. As it can be seen, the FPGA engine achieves up to one to two orders of magnitude speedup when WA is closer to WS. This is because the software implementations benefit from incremental computation by applying an inverse function, whereas the FPGA processes the sliding windows in their entirety. There is also a dip of performance after $WS=k=128$ due to the requirement of an intermediate pass from the merge sorter as illustrated in figure 7b, while still providing competitive performance for the majority of WA-WS combinations, especially for queries like distinct count.
IV-D Experimental results – SWAG engine, with groups
@@
For SWAG with groups, the evaluation also focuses on min./med./max. This operator combines the aggregated results in the form of up to 3 tuples per group representing the minimum, median and maximum with the corresponding group ID in a single run. The logic behind this operator selection is to showcase a variation of algorithms that benefit from the CPU and the FPGA at different degrees. For instance, the sum operator can be calculated incrementally, and a pointer-based solution can potentially have an advantage over the FPGA solution, as it does not have to read entire windows for smaller window updates. On the other hand, min./med./max. is less trivial or rewarding to calculate incrementally on the CPU.
@@
As demonstrated in figure 11, the FPGA advantage of the solution varies for the three tested operators. For the majority of operator, WS and WA combinations, the engine provides a speedup of one to two orders of magnitude. If we ignore the dip after the internal pass is invoked ($\textit{WS}>128$), we can see the upward trend in the figures with respect to the window size, which can be explained by the thoughtful use of the on-chip memory and longer memory transfers, as opposed to the CPU’s cache limitations. An exception is the combination of a high WS and low WA for the sum operator where the FPGA experiences a speedup lower than 1x. This comes from the software advantage of being able to efficiently do incremental updates for such a simple operator as sum, though according to a use case such as with 100% unique group IDs, WS=128, WA=64 the speedup is 164x (not shown).
@@
The studied operators also behave differently under different input distributions at least when compared to the software equivalent. The input of sum and distinct count operators consists of 50% unique group IDs per group, which means every window has around WS/2 unique group IDs. The variations when changing from 100% unique to near 0% unique (single group ID) were relatively limited for these two operators, especially with regard to the FPGA performance. Thus, only the 50% version is shown to conserve space. For example, for WS=16384, WA=8192 the speedup on the distinct count operator is 317x, 420x and 476x for 100%, 50%, and single unique groups respectively. However, this is not the case for the operator min./med./max., hence the overlaid rightmost plots of figure 11 for different distributions. The output size varies according to the distribution, ranging from 1 to 3 tuples per group ID, according to the overlappings. If we select WS=16384, WA=64, the corresponding speedups for min./med./max. change to 59x, 239x and 472x. This is expected as the increased localities from having to operate on fewer elements can benefit software implementations when comparing the overheads of using a co-processor, while noting that the speedup is still impressive.
IV-E Higher-end baselines
Our system-level evaluation focuses on single-thread baselines to avoid introducing more variables. This is inline with reference implementations aiming at algorithmic comparison [19]. Specifically, SWAG exhibits a performance tradeoff between incremental computation and parallelism [20], making the selection of an oracle parallel implementation challenging. For the sake of completeness, a brief comparison is provided with an emulated parallel version on the same platform, and a high-end processor in figure 12. The parallel implementation is emulated by the single-thread version making fork() calls. Then, the median wall-time is divided by the number of “threads”. In this way, multi-processing inefficiencies such as memory pressure are partly reproduced, providing a lower bound for the obtained speedup.
Figure 12: Comparison examples with higher-end baselines, WS=WA=16K
On the same platform, the FPGA speedup ranges from 13x to 148x over the parallel oracle version according to the input and operator. Over 7950X3D, the actual speedup for single-thread is up to 26x, while sum with a single group is slower due to its incremental nature and data localities. When using 32 threads, the performance is similar at up to 1.2x, while noting the same input/operator combination outlier. Note this processor has a more advanced data source (DDR5 vs LPDDR4) and process (5 vs 16 nm). The FPGA power is also notably lower at 3.34W over 45W for this workload (latter influenced by DVFS), highlighting its appropriateness for edge applications, which are prominent with SWAG [1].
IV-F Resource utilisation